Please return to NE43-366

department of electrical engineering and computer science

massachusetts institute of technology

Cambridge, Massachusetts 02139

Application for Approval of Minor in Area II

Please print legibly and in ink. Name of Applicant: Address for reply: Email address: Date: Title or Subject of Minor:

MIT Course No. Credit (U, G, or H)

Grade (if

Prerequisites (Course no.

and title and Units applicable)

Units, +U, G, or H Credit)

1st course name (1st course number)

1st course units 1st course grade

prerequisites

2nd course name (2nd course number)

2nd course units 2nd course grade

prerequisites

3rd course name (3rd course number)

3rd course units 3rd course grade

prerequisites

Non-MIT Course No. and title Where Taken No. of hours

Grade

1st Course name (Course number)

Where taken No. of hours

Grade

2nd Course name (Course number)

Where taken No. of hours

Grade

3rd Course name (Course number)

Where taken No. of hours

Grade

Student’s signature Counselor’s approval Area Chairperson’s approval (initials)